OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x)=\[\frac{ 4 }{ x }\] g(x)=\[\frac{ 4 }{ x }\]

OpenStudy (cwrw238):

f((g(x)) = 4 / (4/x) = 4x / 4 = x now try g((f(x))

OpenStudy (anonymous):

?

OpenStudy (cwrw238):

to find f((g(x)) you replace the x in f(x) by g(x) then simplify